### v3

# GS-Occ3D: Scaling Vision-only Occupancy Reconstruction with Gaussian Splatting

Abstract
Occupancy is crucial for autonomous driving, providing essential geometric priors for perception and planning. However, existing methods predominantly rely on LiDAR-based occupancy annotations, which limits scalability and prevents leveraging vast amounts of potential crowdsourced data for auto-labeling. To address this, we propose GS-Occ3D, a scalable vision-only framework that directly reconstructs occupancy. Vision-only occupancy reconstruction poses significant challenges due to sparse viewpoints, dynamic scene elements, severe occlusions, and long-horizon motion. Existing vision-based methods primarily rely on mesh representation, which suffer from incomplete geometry and additional post-processing, limiting scalability. To overcome these issues, GS-Occ3D optimizes an explicit occupancy representation using an Octree-based Gaussian Surfel formulation, ensuring efficiency and scalability. Additionally, we decompose scenes into static background, ground, and dynamic objects, enabling tailored modeling strategies: (1) Ground is explicitly reconstructed as a dominant structural element, significantly improving large-area consistency; (2) Dynamic vehicles are separately modeled to better capture motion-related occupancy patterns. Extensive experiments on the Waymo dataset demonstrate that GS-Occ3D achieves state-of-the-art geometry reconstruction results. By curating vision-only binary occupancy labels from diverse urban scenes, we show their effectiveness for downstream occupancy models on Occ3D-Waymo and superior zero-shot generalization on Occ3D-nuScenes. It highlights the potential of large-scale vision-based occupancy reconstruction as a new paradigm for scalable auto-labeling. Project Page.
footnotetext: Equal contribution. Listing order is random. Corresponding author.
1 Introduction
Figure 1: Overview of occupancy reconstruction pipelines. a) Existing methods predominantly rely on LiDAR-based occupancy annotations, requiring costly specialized surveying vehicles, which significantly limits scalability. b) In contrast, GS-Occ3D introduces a scalable, vision-only occupancy reconstruction framework that effectively harnesses abundant crowdsourced data from consumer-grade vehicles for auto-labeling. Our approach enables affordable and scalable curation of high-quality occupancy labels. c) We present an overlay of the binary prediction (pink) and the Occ3D-Waymo validation GT (other colors represent semantics), solely to visualize areas where the predictions are incomplete. Comparing models trained with two types of labels, we achieve generally comparable or even better generalization geometry results in certain setups.
Goal: Vision-only reconstruction. Current occupancy reconstruction methods primarily depend on LiDAR-based annotations [30, 104]. As shown in Fig.˜1, these approaches face inherent scalability limitations due to the high costs associated with data collection and annotation. In contrast, vision-centric frameworks offer a promising alternative by leveraging large-scale crowdsourced data for self-supervised auto-labeling, eliminating the reliance on expensive LiDAR sensors [71, 54, 81, 41, 47]. Despite its potential, vision-only occupancy reconstruction remains a challenging task due to the sparsity of viewpoints, severe occlusions, dynamic scene elements, and long-horizon motion. For example, a typical scene in the Waymo dataset contains approximately 1,000 frames captured from only five cameras, resulting in limited co-visible regions across frames and incomplete scene observations.
The success of scalable occupancy reconstruction hinges on an effective scene representation. Recent advances in scene representation learning, including Neural Radiance Fields (NeRF) [55, 1, 56, 73, 70, 86, 92] and 3D Gaussian Splatting (3DGS) [33, 45, 48, 101, 32, 10, 89, 34, 87, 107, 106, 11], have shown remarkable rendering and novel view synthesis capabilities. Among these, 3DGS stands out for its efficiency and speed, but existing methods remain optimized for rendering quality rather than geometric accuracy, making them less suitable for occupancy reconstruction.
Limitations of prior methods. (1) Prior works focusing on geometry reconstruction [9, 16, 21, 96, 27] primarily target indoor, object-centric, or simple outdoor scenes. Extending these techniques to large-scale driving scenes leads to substantial accuracy degradation, particularly in weakly textured regions and along long-horizon, high-speed trajectories. (2) Existing methods that attempt street scene geometry reconstruction [64, 51, 22, 15] typically rely on mesh-based representations, which introduce overly smooth surfaces (in implicit methods) or fragmented, hole-ridden reconstruction (in GS-based methods). These issues necessitate extensive post-processing, hindering their scalability for efficient vision-only label curation. (3) Last but not least, most prior approaches [51, 22, 64, 99, 27, 75, 95, 27, 96] also focus solely on static scenes, failing to capture the dynamic aspects of real-world driving environments. This lack of dynamic object modeling reduces their applicability to downstream tasks [98, 67, 46, 105].
To summarize, scalable vision-only occupancy reconstruction is constrained by three major challenges: (1) Limited geometric priors from sparse views, making accurate reconstruction hard. (2) Degradation over long trajectories, leading to inconsistent and incomplete geometry. (3) Dynamic occlusions from moving objects, creating challenges in accurately modeling interactions in complex scenes.
To overcome these challenges, we introduce GS-Occ3D, a vision-only framework for scalable occupancy label curation that supports large-scale auto-labeling. Our method strategically decomposes driving environments into three geometrically distinct components: (1) Static backgrounds are modeled with an octree-based hierarchical surfel representation for multi-scale fidelity. (2) Ground surfaces are explicitly reconstructed as a dominant structural element to enhance large-area consistency. (3) Dynamic objects are processed separately to better capture motion-related occupancy patterns and reduce occlusion artifacts.
By utilizing this tailored representation, we ensure high spatial-temporal consistency and preserve multi-scale geometric fidelity across long-horizon sequences. With this approach, we reconstruct the entire Waymo dataset to generate vision-only point clouds, which are then processed through a pipeline consisting of frame-wise division, multi-frame aggregation and voxelization. This enables the curation of vision-only occupancy labels, facilitating the training of state-of-the-art downstream occupancy models.
Overall, our contributions are summarized as follows:
  • We introduce a scalable pipeline for vision-only occupancy label curation, eliminating reliance on LiDAR while empowering downstream perception models.
  • Our method effectively reconstructs the ground, background, and dynamic objects from panoramic street views along long trajectories. We outperform existing methods, even surpassing LiDAR-supervised baselines.
  • We are the first to reconstruct the full Waymo dataset using a vision-only approach. We show the effectiveness of our labels for downstream occupancy models on Occ3D-Waymo and superior zero-shot generalization on Occ3D-nuScenes. This highlights the scalability of our approach for large-scale autonomous driving applications.
2 Related Works
Large-scale Scene Reconstruction. With the rapid advancement of NeRF [55, 1, 56, 92, 97] and 3DGS [33], vision-based methods have revolutionized large-scale scene reconstruction. NeRF-based methods segment scenes [70], use grid-based ray association [73], or integrate grids without decomposition [86]. GS-based approaches employ data partitioning for aerial training [45, 48] or ensure global inference consistency [10]. For static street scenes, Hierarchical-GS [34] introduces Level-of-Detail (LOD), UC-GS [101] refines car-view details via cross-view uncertainty, and Horizon-GS [32] unifies aerial-street reconstruction with a coarse-to-fine LOD strategy.
For dynamic scenes, some works [74, 88] model all elements within a single field, while some works [58, 91, 84, 72] decompose scenes into foreground and background for better motion handling. Recent GS-based works [87, 108, 107, 106, 11, 25, 65, 38, 77, 85] further improve the fidelity and efficiency. More recently, some works [36, 53, 93, 23, 90, 57] integrate diffusion models [26, 17] for diverse driving scene generation. However, they rely on LiDAR or prioritize rendering over geometry. We address this gap with a vision-only reconstruction method focused on geometry.
Vision-only Geometry Reconstruction. Vision-only geometry reconstruction in urban scenes is challenging due to dynamic objects, sparse views, occlusion and long-horizon motion. NeRF-based methods like DNMP [51] use neural mesh primitives, while StreetSurf [22] incorporates monocular cues (e.g., depth, normal) from pretrained models. However, their reliance on complete meshes or full-volume processing limits scalability. GS-based methods advance object-centric and bounded scene reconstruction [9, 16, 21, 96, 99, 27, 75, 95], but scaling to street scenes faces issues like uneven point distribution, holes, and floaters, making them unsuitable for downstream tasks. Recent grid-based improvements [64, 15] enhance mesh quality but struggle in weak-texture regions. In contrast, we reconstruct both ground and dynamic geometry, advancing scalable vision-only occupancy reconstruction and enabling seamless label curation for downstream tasks.
3D Occupancy Prediction. 3D occupancy prediction estimates the occupancy of each voxel in 3D space. With the rise of multiple benchmarks [71, 81, 41, 47, 82] based on large-scale datasets [69, 5, 2, 44, 8], this task has gained traction, especially for dynamic street-view applications.
Most methods rely on costly LiDAR-based occupancy labels. LiDAR-based methods [66, 63, 37] complete scene occupancy from sparse LiDAR inputs, while camera-based methods predict 3D occupancy by extracting features from 2D images [6, 40, 39, 43, 102, 28, 42, 94, 59, 50, 83, 61, 12]. Monocular methods infer 3D structure from a single image via 2D-to-3D backprojection [6] or depth-aware cross-attention [40], while multi-view approaches generate 3D volumetric features from multiple camera views [39, 43, 102, 28, 42, 94, 59, 50, 83]. Some works use GS to transform 2D images into dense gaussian representations [30, 104]. Despite strong results, they heavily depend on LiDAR-based annotations, which is costly and time-consuming. Recent vision approaches reduce reliance on LiDAR by using volume rendering with 2D supervision [29, 60, 100, 61, 103, 4, 7, 49] or pretrained vision-language models [109, 3, 31]. Instead, we focus on leveraging vision-only 3D geometry labels, which is cheaper, more scalable and efficient for downstream occupancy models.
3 Methods
Fig.˜2 illustrates the overview of our framework. We first generate a sparse point cloud and ground surfels from panoramic street views captured along long trajectories. For scalable vision-only geometry reconstruction, we use an octree-based Gaussian Surfel representation integrating ground, background, and dynamic objects. Our label curation pipeline refines the vision-only point cloud through frame-wise division and multi-frame aggregation, increasing density, especially for dynamic objects. Ray-casting then resolves occlusions for accurate voxel occupancy labeling. The resulting vision-only labels can supervise downstream occupancy models, improving their generalization and geometric reasoning capabilities.
3.1 Scalable Vision-only Geometry Reconstruction
Figure 2: Overview of the GS-Occ3D. Left: Panoramic street views captured along long trajectories are used to generate a sparse point cloud and ground surfels as initialization. We adopt Octree-based Gaussian Surfel representation that integrates ground, background, and dynamic objects to achieve scalable vision-only geometry reconstruction. Here, we present an uphill scene with colors indicating height. Middle: Given the vision-only point cloud, our label curation pipeline applies frame-wise division and multi-frame aggregation to define appropriate perception ranges per frame, while increasing point cloud density, especially for dynamic objects with incomplete observations. Ray-casting is then applied to each frame to determine voxel occupancy, explicitly handling occlusions from the camera’s viewpoint. Right: The resulting vision-only labels can be used to train downstream occupancy models, enabling these models to generalize to unseen scenes with geometric reasoning capability. Pink indicates the binary voxel, while other colors represent Occ3D labels.
Preprocessing. In the absence of geometric priors, we exclusively use detector-free SfM [68, 24] and ground gaussians detailed later to generate sparse point clouds as the initial scene representation. We also use an off-the-shelf segmentation model [14] for necessary decoupling.
Octree-based Gaussian Surfel. To address the inherent lack of geometric priors in sparse views conditions, we adopt an octree-based gaussian surfel inspired by [62], using initial sparse point clouds as the scene skeleton to maximize the use of geometric information from the preprocessing stage. This structure enables hierarchical spatial partitioning and flexible adjustment during training, ensuring efficient, accurate, and scalable geometry reconstruction.
The dynamic octree structure adapts during training, expanding or contracting to match scene density and complexity. Each sparse voxel can generate up to $m$ Gaussian primitives, confined to a small region centered at the voxel. These primitives serve as localized surfel representations, effectively approximating local surface geometry.
The voxel resolution adapts across different octree levels. Coarser levels feature lower spatial density, efficiently modeling global structures like walls and roads, while finer ones capture high-frequency details such as vegetation, buildings and object boundaries. The number of octree levels $K$ is determined by the observed distribution of distances between camera centers and input sparse point cloud, following a strategy similar to  [62]. It is computed as:
$K=\left\lfloor\log_{2}\left(\frac{d_{\max}}{d_{\min}}\right)\right\rceil+1.$ (1)
where $d_{\min}$ and $d_{\max}$ are the smallest and largest distances between camera centers and SfM points. This ensures the octree provides sufficient granularity to capture both near- and far-field structures within the scene.
Once $K$ is determined, we initialize voxel centers at each octree level. Starting with a base voxel size $\epsilon$ at the coarsest level, the voxel centers at level $L$ are computed as:
$\mathbf{V}_{L}=\left\{\left\lfloor\frac{\mathbf{P}}{\epsilon/2^{L}}\right\rceil\cdot\frac{\epsilon}{2^{L}}\right\}.$ (2)
Here, $\mathbf{P}$ denotes the 3D coordinates of the sparse point cloud. This hierarchical quantization keeps the voxel centers spatially aligned across different levels, which is essential for efficient hierarchical geometry representation. Finally, after constructing the octree and initializing voxel centers, we assign $m$ Gaussian surfels to each voxel.
We use an adaptive anchor control strategy inspired by [52, 62] to manage surfels. To enhance geometry completeness across scales, we apply cumulative LOD levels instead of a single one, to capture both coarse scene coverage and fine geometric details for high-fidelity reconstruction.
Ground Reconstruction. Existing reconstruction methods, despite incorporating geometric constraints, struggle with weakly-textured ground in sparse-view scenes, highlighting the need for a dedicated ground modeling approach.
Assuming the road is parallel to the cameras, we initialize Ground Gaussian surfels by projecting camera poses onto the xy-plane, inspired by [19]. To handle elevation, each surfel’s z-coordinate is adjusted using the nearest camera pose with a fixed height offset, while its orientation inherits the nearest camera’s rotation. This initialization aligns surfels with the road slope, enabling adaptation to various terrains as shown in Fig.˜3. Finally, planar regularization encourages smooth planar structures, further improving ground geometry as shown in Fig.˜8.
Dynamic Reconstruction. We assume each dynamic vehicle is associated with a 3D bounding box predicted from RGB images and represented by points with tracked poses $\mathbf{R}_{t}$ and $\mathbf{t}_{t}$. We initialize dynamic vehicles using vision-based 3D object tracking method [20]. A fixed number of points are then sampled within each box.
To mitigate noise in the initial poses, we further refine them with learnable corrections, following  [87]:
$\mathbf{R}_{t}^{\prime}=\mathbf{R}_{t}\Delta\mathbf{R}_{t},\quad\mathbf{t}_{t}^{\prime}=\mathbf{t}_{t}+\Delta\mathbf{t}_{t},$ (3)
The learnable translation $\Delta\mathbf{t}_{t}$ and rotation $\Delta\mathbf{R}_{t}$ enable efficient training via direct gradient computation.
Loss Function. The total loss combines weighted RGB, geometry, object, ground, and sky components:
$\begin{split}L=&L_{rgb}+\lambda_{geo}L_{geo}+\lambda_{obj}L_{obj}\\ &+\lambda_{road}L_{road}+\lambda_{sky}L_{sky},\end{split}$ (4)
where $\lambda_{geo}$, $\lambda_{obj}$, $\lambda_{road}$, and $\lambda_{sky}$ denote the corresponding weights. The geometry loss $L_{geo}$ consists of three terms: surfel regularization, depth distortion, and depth-normal consistency, formulated as:
$L_{geo}=\lambda_{s}L_{s}+\lambda_{d}L_{d}+\lambda_{n}L_{n}.$ (5)
Here, the RGB loss $L_{rgb}$ integrates L1 and D-SSIM losses to supervise RGB reconstruction, following [33]. The surfel regularization loss $L_{s}$ flattens Gaussians into surfels. The depth distortion loss $L_{d}$ and the normal depth consistency loss $L_{n}$ encourage surfels to better conform to the geometry of the underlying scene, following the geometry constraints of [27, 96]. Object loss $L_{obj}$ applies an entropy loss to the object opacity map, encouraging clearer decoupling between foreground and background. Road smoothness loss $L_{road}$ preserves flatness by regularizing height variations between neighboring surfels. Sky loss $L_{sky}$ applies a binary cross-entropy loss to the rendered opacity.
3.2 Occupancy Labels Curation
With scalable geometry reconstruction, it becomes feasible to curate vision-only 3D occupancy labels. However, the relatively sparse point cloud produced during geometry reconstruction makes it hard to obtain dense voxel representations, which leads to the sparsity problem. As the point cloud is densified, identifying occluded and invisible voxels from the camera views becomes essential. To address these issues, we employ frame-wise division and multi-frame aggregation to define per-frame perception ranges for vision-only point clouds and increase point-cloud density, particularly for dynamic objects with incomplete observations. Then we can operate ray-casting to determine the occupancy status of every voxel, explicitly handling occlusion.
Frame-Wise Division. Unlike incremental LiDAR sequences, the reconstructed point cloud covers the entire scene in batches, which requires frame-wise division. To achieve this, we define a perception range centered on the camera pose, approximating the typical sensing range of LiDAR. Within this range, we uniformly sample points to form a single-sweep point cloud, ensuring that the number of points is consistent with that of a real LiDAR sweep.
Multi-frame Aggregation. The reconstructed point cloud is relatively sparse, particularly for dynamic objects. To mitigate this, we aggregate points belonging to dynamic objects across frames, increasing their density.
Directly merging points across frames can cause smearing or distortion in dynamic objects like vehicles, so it is necessary to process them separately. Since our pipeline explicitly decouples static and dynamic components, we can extract optimized points within tracked bounding boxes without additional segmentation. These points are transformed from the sensor coordinate system to the box coordinate system, following a process similar to [71]. By concatenating these transformed points across frames, we effectively densify the point clouds for dynamic objects.
For static scenes, frame aggregation is unnecessary, as the static point cloud sequence is directly obtained by slicing the reconstructed scene. Unlike LiDAR-based labels, where frames provide complementary observations, static frames in our approach are inherently complete. After placing the densified dynamic points back into their corresponding bounding boxes in each frame, we fuse the static scene with the aggregated dynamic objects in the current frame, producing a dense single-frame point cloud.
Voxelization. To generate a 3D occupancy grid from aggregated point clouds, a straightforward approach is to mark voxels containing points as occupied and others as free. However, due to the limited camera field of view, some occupied voxels are only partially observed or entirely occluded from the camera’s perspective, which can lead to incorrect labeling as free. This ambiguity can confuse downstream models during training, making it necessary to distinguish between free and unobserved voxels. Inspired by [71], we use ray-casting to determine voxel visibility. We cast a ray from the camera to each occupied voxel. Only the first occupied voxel encountered along any ray is labeled as observed, while all others are considered unobserved.
4 Experiments
Figure 3: Visualization of Geometry Reconstruction on Waymo. The color represents CD with respect to LiDAR, ranging from blue (lower CD) to red (higher CD). We exhibit improved reconstruction fidelity in weakly-textured regions compared to other methods, while maintaining structural completeness comparable to LiDAR point cloud, even in the absence of geometric priors.
4.1 Geometry Reconstruction Experimental Setup
Most geometry reconstruction methods focus on static scenes. For fairness, we compare our reconstructed static components with SOTA static scene methods. Following [88], we evaluate on the Waymo Static-32 split and select the most scalable method with high geometric accuracy to reconstruct the full Waymo dataset and curate labels.
Datasets. To evaluate performance and scalability in large-scale open scenes, we use the Waymo dataset [69]. We use all five camera views and all frames in each scene, resulting in about 1,000 images per scene. LiDAR point clouds are used as reference to assess geometric accuracy.
Baselines. For geometry reconstruction, we evaluate our method against SOTA implicit methods (NeuS [78], F2-NeRF [79], StreetSurf [22]) and GS-based methods (2DGS [27], PGSR [9], GVKF [64]). All explicit baselines use the same input point clouds for initialization.
Metrics. We evaluate reconstruction quality across geometry, rendering, and efficiency. For geometry, we measure the accuracy of both point clouds and meshes using Chamfer Distance (CD). For rendering quality, we report peak signal-to-noise ratio (PSNR). For efficiency, we record storage requirements, GPU memory usage, and training time.
4.2 Geometry Reconstruction Result Analysis
Figure 4: Details of Geometry Reconstruction. We present the detailed geometry of the red-boxed area in Fig.˜3, achieving results that are both comparable and complementary to LiDAR. The first row is uphill, while the second is downhill followed by uphill.
SOTA Geometry Reconstruction. Footnote˜3 shows that we achieve SOTA geometry reconstruction, surpassing methods that use monocular cues or LiDAR-input or supervised like StreetSurf and NeuS, which demonstrates the reliability of geometry reconstruction. Besides geometry, we also maintain competitive rendering quality and training efficiency. This is partly due to the octree-based representation, which is both memory efficient and faster to train. Although we reconstruct denser point clouds, the structured octree helps accelerate geometry convergence while ensuring high rendering and geometric accuracy. Under sparse-view conditions, the octree-based gaussian surfel with LoD preserves geometric fidelity through multi-scale consistency.
Comparable and Complementary to LiDAR. As shown in  Fig.˜3 and  Fig.˜4, our point cloud is globally comparable to LiDAR. In certain regions, such as tall buildings and thin poles, we achieve even higher quality, capturing finer details. Additionally, it provides more detailed textures and reconstructs areas beyond LiDAR’s limited range, making it reliable for downstream tasks. As observed in the comparison, our reasonable reconstruction beyond the LiDAR coverage is also counted as regions with high CD.
Method Geometry Rendering Efficiency
CD $\downarrow$ PSNR $\uparrow$ MB $\downarrow$ GB $\downarrow$ Time $\downarrow$
Implicit NeuS11footnotemark: 1 [78] 0.76 13.24 170 31 5.0h
$\text{F}^{2}$-NeRF [79] 886.77 24.70 130 24 0.8h
StreetSurf [22] 1.02 27.12 540 22 1.5h
StreetSurf22footnotemark: 2 [22] 0.90 26.85 245 21 1.5h
Explicit PGSR33footnotemark: 3 [9] 2.90 22.61 78 4 1.5h
2DGS33footnotemark: 3 [27] 1.23 25.60 83 15 1.0h
GVKF33footnotemark: 3 [64] 0.82 25.87 65 24 2.0h
Ours 0.56 26.89 80 10 0.8h
Table 1: Performance of implicit and explicit geometry reconstruction methods on the Waymo Static-32 Split. NeuS11footnotemark: 1 uses 1 dense and 4 sparse LiDARs, StreetSurf22footnotemark: 2 uses 4 sparse LiDARs, and all other methods are vision-only.33footnotemark: 3 indicates using our ground gaussians. MB indicates storage size, GB indicates GPU memory, and Time indicates training time.
Figure 5: Qualitative Results of Our Curated Labels. We achieve globally comparable geometry to Occ3D, ensuring reliable supervision for occupancy model training without priors.
4.3 3D Occupancy Prediction Experimental Setup
Through reconstructing the full Waymo dataset, we find vision-only methods fail to handle ego-static scenes. So we exclude these unreliable parts from the labels for training occupancy models. The experiments compare downstream occupancy models trained with the labels from Occ3D [71] and our labels. We evaluate the generalization performance on the validation set of Occ3D-Waymo and Occ3D-nuScenes and the fitting performance on the training set.
Datasets. Occ3D-Waymo is among the most diverse and comprehensively labeled open-source 3D occupancy datasets [71], which contains 798 training scenes and 202 validation scenes, covering approximately 200,000 frames in total. For fairness, we exclude ego-static scenes that cannot be reliably handled by vision-only reconstruction, resulting in 637 training scenes and 165 validation scenes. We also use original Occ3D-nuScenes validation set including 150 scenes for evaluating zero-shot generalization.
Baselines. We compare our labels with LiDAR-based labels Occ3D [71] using SOTA occupancy model CVT-Occ [94] which leverages geometric correspondences of 3D voxels over time to improve occupancy prediction accuracy.
Metrics. To evaluate the geometric accuracy and reliability of our labels, we use the Intersection over Union (IoU) to assess geometry performance of 3D occupancy prediction. For fair comparison, the evaluation only considers voxels within the visible region of the camera views.
4.4 3D Occupancy Prediction Result Analysis
Train Labels Eval Labels IoU $\uparrow$ F1 $\uparrow$ Prec. $\uparrow$ Rec. $\uparrow$
Ours (Waymo) Occ3D-Val(Waymo) 44.7 61.8 58.2 65.9
Occ3D (Waymo) 57.4 73.0 62.9 87.0
Ours (Waymo) Occ3D-Val(nuScenes) 33.4 50.1 62.5 41.8
Occ3D (Waymo) 31.4 47.8 38.8 62.1
Ours (Waymo) Ours-Val(Waymo) 46.8 63.8 54.6 76.6
Occ3D (Waymo) 41.1 58.3 46.7 77.6
Ours (Waymo) Ours-Train(Waymo) 50.6 67.2 54.3 88.1
Occ3D-Train(Waymo) 48.3 65.1 60.1 71.0
Table 2: Generalization and Fitting Results on the Occ3D Dataset for the SOTA occupancy model CVT-Occ [94] under different training and evaluation label combinations.
Figure 6: Generalization Results on the Occ3D-Waymo Validation Set. We evaluate the SOTA occupancy model CVT-Occ [94] trained with our labels and Occ3D, achieving reasonable and overall comparable results.
Label Curation Results.  Fig.˜5 visualizes the comparison between our labels and Occ3D. Without geometric priors and initializing with sparse point clouds, we still achieve globally comparable geometry to the LiDAR-based labels. This ensures reliable geometry for downstream models.
Comparable and Superior Zero-shot Occupancy Generalization Results.  Fig.˜6 and  Tab.˜2 show the generalization results of CVT-Occ [94] on the validation set. We demonstrate the ability to reconstruct generalizable geometry, achieving reasonable and overall comparable results on Occ3D-Waymo. Despite inherent camera limitations (e.g., Waymo’s forward-facing views versus LiDAR’s 360-degree coverage), we achieve slightly lower performance on the Occ3D-Val(Waymo) but still within a reasonable range. While both we and Occ3D perform better on our respective validation sets, we notably demonstrate superior zero-shot generalization on nuScenes with diverse camera settings, yielding more complete geometry in textured and distant regions. This is particularly impressive given that Occ3D relies on high-end LiDAR data. We rely purely on camera inputs yet delivers comparable or even better generalization performance in certain setups. This highlights not only the scalability of vision-only methods, but also their potential to match or surpass LiDAR-based baselines in real world.
Occupancy Fitting Results. As shown in the last two rows of Tab.˜2, the performance gap evaluated on Occ3D-Train and Ours-Train is small, and the precision on the Occ3D-Train and Occ3D-Val remains similar. This indicates that our labels provide learnable geometry, enabling the model to capture meaningful geometry information.
Figure 7: More Visualization. Up: Richer semantic labels. Down: superior generalization on Occ3D-nuScenes. Pink indicates binary prediction, others show errors.
Advantages over LiDAR. (1) Wider coverage: Vision-only reconstruction mitigates LiDAR’s limited spatial range, excelling in large-scale areas and tall structures like high-rise buildings. (2) Superior zero-shot generalization: Models trained with our labels can predict a wider range of geometries and generalize better to unseen scenes than LiDAR-based ones. (3) Cheap and Rich Semantics: Unlike LiDAR semantics which require labor-intensive annotation or costly multi-sensor fusion, RGB images inherently capture color, texture, and object-class cues. Leveraging 2D segmentation models like  [14], we produce a richer 3D category set (66 vs. Occ3D’s 16), identifying objects LiDAR misses yet vital for driving like motorcycles, lane markings and crosswalks. (Fig.˜7). (4) Greater potential in adverse weather: Although adverse weather impacts both sensors, vision can use rich textures and semantics and learned priors from large-scale data, to reconstruct degraded scenes more effectively. From Fig.˜7, we even generalize better than LiDAR in rainy scenes.
4.5 Ablation Studies
Figure 8: Ablation Results of Ground Gaussians. We show the effectiveness of our ground gaussians. Colors indicate height, ranging from blue to red.
Views Method Geometry Rendering
CD-Pcd $\downarrow$ CD-Mesh $\downarrow$ PSNR $\uparrow$
5Cam PGSR [9] 3.63 4.41 19.18
2DGS [27] 1.25 2.14 23.42
GVKF [64] 0.82 1.22 25.87
Ours 0.56 - 26.89
3Cam PGSR [9] 2.90 3.03 22.61
2DGS [27] 1.23 1.85 25.60
GVKF [64] 0.87 1.02 26.22
Ours 0.66 - 26.96
Table 3: Ablation Studies on Waymo Static-32 Split. We evaluate varying camera counts and representations. All methods use our Ground Gaussians for fairness. Both chamfer distance of the point cloud and mesh are measured against LiDAR.
Input Camera Views. As shown in Tab.˜3, our method improves reconstruction with a 5-camera input over 3. In contrast, the performance degradation in other methods is primarily caused by the geometric and rendering ambiguities from forward-facing multi-view inputs. However, our octree-based representation with LoD maintains structural fidelity and consistency across both global and local scales, enabling effective exploitation of multi-view observations. This makes our method well-suited for reconstructing diverse crowdsourced data and generating high-quality label.
Choice on Point Cloud or Mesh. Tab.˜3 depicts that, for GS-based methods, directly representing geometry as a point cloud is preferable to mesh conversion. Mesh reconstruction introduces post-processing that transforms gaussians into surfaces, leading to information loss due to incomplete observations and the inherent limitations of meshing algorithms. This process often results in holes, sky enclosure artifacts, where the sky is misrepresented as a surface wrapping around the scene, and other errors that require further post-processing, which reduces scalability.
Effectiveness of Our Ground Gaussians. Fig.˜8 shows the effectiveness in capturing geometry of weakly-textured areas. Without separate handling of the ground, holes and abnormal protrusions can occur, distorting the geometry.
5 Conclusion
GS-Occ3D is a scalable vision-only occupancy reconstruction framework that uses crowdsourced data from consumer vehicles for auto-labeling. It enables cost-efficient and scalable occupancy label curation. First, we adopt an Octree-based Gaussian Surfel formulation to reconstruct geometry for ground, background, and dynamic objects. Our vision-only reconstruction achieves SOTA geometric results. We then reconstruct the entire Waymo dataset, which covers diverse scenes. This enables us to validate the effectiveness of our labels for downstream occupancy models on Occ3D-Waymo and demonstrate superior zero-shot generalization on Occ3D-nuScenes. Our results highlight the potential of large-scale vision-based occupancy reconstruction as a new paradigm for scalable auto-labeling. We hope this work inspires further research into using advanced reconstruction to empower vision-only label curation for downstream tasks.
Limitations. (1) Cameras inherently provide only front and side views, lacking rear view coverage, which leads to unavoidable information loss. (2) Performance degrades at night or with exposure issues, which reduce the effective visual range. (3) Vision methods often fail in ego-static scene. While methods pretrained on Waymo LiDAR [80, 35, 18] are incompatible with our vision-only setting, recent works [76, 13] may offer solutions. (4) We focus on geometry reconstruction, providing results solely on geometry generalization. Future work will aim to enhance both semantic and geometry reconstruction, particularly under challenging weather and lighting conditions, to achieve greater robustness in real-world applications.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (2022YFB4701900), and the Tsinghua University - Mercedes Benz Joint Institute for Sustainable Mobility.